OpenStudy (anonymous):

multiply and check (4x-5x-3)(1+2x^2)

Parth (parthkohli):

\[(4x - 5x - 3) (1 + 2x^2) = 1(4x - 5x - 3) + 2x^2(4x - 5x - 3) =\cdots\]

OpenStudy (anonymous):

Use Distributive Law

OpenStudy (anonymous):

-2x^(3)-6x^(2)-x-3

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

he gave you the work. i gave you the answer

OpenStudy (anonymous):

lol i dont see the work any where

OpenStudy (mathstudent55):

Multiply every term inside the left parenteses by every term inside the right parentheses. Then collect like terms. But before you do that, I have a question. Is the left side correct? Is it really 4x - 5x - 3? There is no square on either x? If 4x - 5x - 3 is correct, then first simplify it to -x - 3, then multiply it by 1 + 2x^2 the way I wrote above.